# Property Enhanced Instruction Tuning for Multi-task Molecule Generation with Large Language Models

### v2

@@ 5 Conclusion
We propose a novel framework PEIT that aims to enable open-source LLMs to perceive multi-modal features for multi-task molecule generation. For this purpose, PEIT establishes cross-modal connections among molecular structures, textual description, and biochemical properties through multi-modal representation alignment. Through template filling, PEIT can help synthesizing diverse task-specific instruction data for LLMs. We further introduce a new multi-constraint molecule generation task that requires generating novel molecules meeting multiple property constraints. Experiments show that PEIT achieves promising performances on molecule captioning, text-based molecule generation, and property-related tasks compared with various biomolecular models and LLMs.
@@ Appendix A Template Filling
Appendix B Details of Classification Tasks
Following SPMM Chang and Ye (2024) , we adopt four commonly-used binary classification tasks to evaluate the performance of PEIT-GEN, including BBBP, BACE, Clintox, and SIDER dataset. The BBBP dataset contains 2,050 molecular samples and aims to predict whether these molecules can cross the blood-brain barrier. The BACE dataset includes 1,513 molecular samples and is used to predict whether a molecule can inhibit the activity of the BACE1 enzyme. The Clintox dataset contains 1,478 molecular samples and is primarily used to predict the toxicity of compounds. The SIDER dataset consists of 1,427 drug samples and is used to predict whether a drug will cause specific side effects. Specifically, we use scaffold splitting and each dataset is divided into a training set, validation set, and test set in a ratio of 8:1:1, respectively.
@@ Appendix C Details of Baselines
GROVERlarge Rong et al. (2020) leverages multiple self-supervised learning tasks to learn universal representations of atoms and bonds in molecular structures, significantly enhancing performance in downstream tasks such as molecular property prediction and drug discovery.
@@ Appendix D Evaluation Metrics
We evaluated the quality of generated text using BLEU Papineni et al. (2002), METEOR Banerjee and Lavie (2005), and ROUGE scores. These metrics evaluate the similarity between generated texts and reference descriptions, effectively quantifying the accuracy and diversity of the generated descriptions. For the text-based molecule generation task, we further use molecular fingerprints (FTS) Cereto-Massagué et al. (2015) and validity measures to assess molecular similarity and validity, including Validity, Levenshtein Levenshtein (1966), MACCS FTS, Morgan FTS, and RDKit FTS Landrum et al. (2013). For the task of molecular property prediction, we chose to use the commonly used RMSE to measure the difference between the predicted values and the molecular property values calculated by RDKit for comparison, for the experiments on MoleculeNet, we use AUC-ROC to evaluate the classification accuracy for classification tasks. In the case of multi-constraint molecule generation, in addition to RMSE, we also employed R2 to assess the accuracy of the generated molecules.
| Module | Parameters |
| Encoder | 440M |
| Encoder momentum cache | 440M |
| Projection head | 1.5M |
| ITM head | 0.6M |
| Property prediction module | 1M |
| Text prediction module | 1M |
| Total | 884.1M |
@@
Appendix E Parameters Analysis
We conduct a detailed analysis of the parameter counts across different modules in PEIT-GEN. As shown in Table 6, the encoders for three modalities are responsible for learning representations of different data types and facilitating the effective fusion of multi-modal information, accounting for approximately 99% of the total size. The remaining modules, such as the projection head, ITM head, property prediction as well as text prediction modules, collectively account for 1% of the total parameter count.
Appendix F Molecular Property Prediction
Following SPMM Chang and Ye (2024), we further compare PEIT-GEN with SPMM on external dataset. The comparison result on molecular property prediction is shown in Table 7. Specifically, we randomly sample 1,000 molecules from the ZINC dataset which are not included in the training set. Compared to SPMM that is specifically designed for property prediction, PEIT-GEN achieves comparable performance while using only one-third of the data size across three modalities. We found that PEIT-GEN outperformed SPMM in terms of RMSE, while SPMM was slightly ahead by 0.11% on R2 metric. These results demonstrate that PEIT-GEN can generate high-quality biochemical properties of molecules, highlighting the critical role of high-quality multi-modal data in advancing molecular understanding tasks.
Appendix G More Analysis on SFT Steps
The impact of SFT steps for text-based molecule generation and single MolLogP property prediction are shown in Figure 6 and Figure 7, respectively. Similar to two other tasks, we receive the best results through approximately 56 epochs of SFT.
Appendix H Case Study

### v4

@@ 5 Conclusion
We propose PEIT, a framework that enables LLMs to perceive multi-modal features for multi-task molecule generation. PEIT aligns molecular structures, textual descriptions, and biochemical properties through multi-modal representation learning. It leverages templates to synthesize diverse, task-specific instruction data for LLMs. We also introduce a challenging multi-constraint molecule generation task, which requires generating novel molecules that satisfy multiple property constraints. Results show that PEIT outperforms various biomolecular models and LLMs on captioning, generation, and property prediction tasks.
@@ Appendix A Template Filling
Appendix B Details of Property Prediction Tasks
Following SPMM Chang and Ye (2024) , we adopt four commonly-used binary property prediction tasks to evaluate the performance of PEIT-GEN, including BBBP, BACE, Clintox, and SIDER dataset. The BBBP dataset contains 2,050 molecular samples and aims to predict whether these molecules can cross the blood-brain barrier. The BACE dataset includes 1,513 molecular samples and is used to predict whether a molecule can inhibit the activity of the BACE1 enzyme. The Clintox dataset contains 1,478 molecular samples and is primarily used to predict the toxicity of compounds. The SIDER dataset consists of 1,427 drug samples and is used to predict whether a drug will cause specific side effects. Specifically, we use scaffold splitting and each dataset is divided into a training set, validation set, and test set in a ratio of 8:1:1, respectively.
@@ Appendix C Details of Baselines
GROVER Rong et al. (2020) leverages multiple self-supervised learning tasks to learn universal representations of atoms and bonds in molecular structures, significantly enhancing performance in downstream tasks such as molecular property prediction and drug discovery.
@@
GPT-3.5 Turbo OpenAI (2023) is an advanced large language model developed by OpenAI, optimized for efficient inference and versatile natural language understanding and generation tasks. Built upon the transformer architecture, GPT-3.5 Turbo demonstrates strong performance across a wide range of NLP benchmarks, including text completion, summarization, translation, and dialogue systems. Its design balances high accuracy with reduced computational cost, making it suitable for scalable real-world applications.
@@
ChemDFM Zhao et al. (2025) is a large language foundation model for the field of chemistry. Trained on 34 billion tokens from chemical literature and 2.7 million instructions, it demonstrates strong capabilities in understanding and reasoning about chemical knowledge. It supports tasks such as molecule recognition, design, property prediction, and reaction analysis, outperforming many open-source large language models.
ChatGLM GLM et al. (2024) is an open-source bilingual large language model optimized for Chinese and English. It supports instruction tuning and multi-round dialogue, making it adaptable to domain-specific tasks such as molecular captioning and property prediction with appropriate prompting.
Galactica Taylor et al. (2022) is a large language model pretrained on scientific texts, including papers, molecules, and protein sequences. Designed to assist scientific reasoning and knowledge retrieval, it supports molecule-related tasks to a limited extent through its exposure to structured scientific data during pretraining.
Vicuna Chiang et al. (2023) is an open-source language model fine-tuned from LLaMA using user-shared conversations. It focuses on improving instruction-following and dialogue capabilities, and can be adapted to domain-specific tasks through fine-tuning, despite lacking scientific domain pretraining.
@@ Appendix D Evaluation Metrics
We evaluated the quality of generated text using BLEU Papineni et al. (2002), METEOR Banerjee and Lavie (2005), and ROUGE scores. These metrics evaluate the similarity between generated texts and reference descriptions, effectively quantifying the accuracy and diversity of the generated descriptions. For the text-based molecule generation task, we further use molecular fingerprints (FTS) Cereto-Massagué et al. (2015) and validity measures to assess molecular similarity and validity, including Validity, Levenshtein Levenshtein (1966), MACCS FTS, Morgan FTS, and RDKit FTS Landrum et al. (2013). For the task of molecular property prediction, we chose to use the commonly used RMSE to measure the difference between the predicted values and the molecular property values calculated by RDKit for comparison, for the experiments on MoleculeNet, we use AUC-ROC to evaluate the accuracy for property prediction tasks. In the case of multi-constraint molecule generation, in addition to RMSE, we also employed R2 to assess the accuracy of the generated molecules.
@@
Appendix E Molecular Property Prediction
Following SPMM Chang and Ye (2024), we further compare PEIT-GEN with SPMM on external dataset. The comparison result on molecular property prediction is shown in Table 7. Specifically, we randomly sample 1,000 molecules from the ZINC dataset which are not included in the training set. Compared to SPMM that is specifically designed for property prediction, PEIT-GEN achieves comparable performance while using only one-third of the data size across three modalities. We found that PEIT-GEN outperformed SPMM in terms of RMSE, while SPMM was slightly ahead by 0.11% on R2 metric. These results demonstrate that PEIT-GEN can generate high-quality biochemical properties of molecules, highlighting the critical role of high-quality multi-modal data in advancing molecular understanding tasks. To further illustrate the predictive performance of PEIT-GEN on the 53 molecular properties, we calculated the relative difference of these properties, as shown in Figure 8.
Appendix F Case Study